# Symmetry Breaking in an Extended O(2) Model

### 29 Dec 2023 v1

@@ Appendix A Connecting to the Clock Model
For non-integer $q$, the $2\pi$-rotational symmetry is broken. Thus, we have to carefully define the domain of the angles $\varphi_{x}$. For example, one could choose $\varphi_{x}\in[-\pi,\pi)$ or $\varphi_{x}\in[0,2\pi)$. These different choices lead to very different phase diagrams as we showed for the $h_{q}\rightarrow\infty$ case in [7]. We prefer the choice $\varphi_{x}\in[0,2\pi)$, simply because the phase diagram gives a more consistent periodic picture as $q$ is increased [^3] See the left panel of Figure 26 in [7] as opposed to the right panel in the same figure.. However, with the choice $\varphi_{x}\in[0,2\pi)$, there is a hard cutoff at the angle $\varphi=0$. When $h_{q}=\infty$, this is not a problem, but at finite $h_{q}$, this cutoff skews the distribution of angles severely enough that the model at finite $h_{q}$ does not smoothly connect to the model at infinite $h_{q}$ when $h_{q}\rightarrow\infty$. To fix this, we need to slightly shift the angle domain such that $\varphi_{x}\in[-\varepsilon,2\pi-\varepsilon)$. To match the clock model (i.e. $h_{q}=\infty$ case), one needs an $\varepsilon$ that varies with $q$ and satisfies the condition $0<\varepsilon<2\pi(1-\lfloor q\rfloor/q)$. In this work, we choose
@@
When $q$ is noninteger, one may find that the Extended O(2) model with $h_{q}\rightarrow\infty$ and the Extended $q$-state clock model do not agree when $\beta\rightarrow 0$. For non-integer $q$, the $\mathbb{Z}_{q}$ symmetry is explicitly broken, and for example, the energy density of the Extended $q$-state clock model does not go to zero in the limit $\beta\rightarrow 0$ [7]. In contrast, the energy density in the Extended O(2) model does go to zero in the limit $\beta\rightarrow 0$ for all values of $h_{q}$. The discrepancy can be understood by considering the Boltzmann factor of the Extended O(2) model, which has the form

### 26 Feb 2024 v2

@@ Appendix A Connecting to the Clock Model
For noninteger $q$, the $2\pi$-rotational symmetry is broken. Thus, we have to carefully define the domain of the angles $\varphi_{x}$. For example, one could choose $\varphi_{x}\in[-\pi,\pi)$ or $\varphi_{x}\in[0,2\pi)$. These different choices lead to very different phase diagrams as we showed for the $h_{q}\rightarrow\infty$ case in [7]. We prefer the choice $\varphi_{x}\in[0,2\pi)$, simply because the phase diagram gives a more consistent periodic picture as $q$ is increased [^3] See the left panel of Figure 26 in [7] as opposed to the right panel in the same figure.. However, with the choice $\varphi_{x}\in[0,2\pi)$, there is a hard cutoff at the angle $\varphi=0$. When $h_{q}=\infty$, this is not a problem, but at finite $h_{q}$, this cutoff skews the distribution of angles severely enough that the model at finite $h_{q}$ does not smoothly connect to the model at infinite $h_{q}$ when $h_{q}\rightarrow\infty$. To fix this, we need to slightly shift the angle domain such that $\varphi_{x}\in[-\varepsilon,2\pi-\varepsilon)$. To match the clock model (i.e. $h_{q}=\infty$ case), one needs an $\varepsilon$ that varies with $q$ and satisfies the condition $0<\varepsilon<2\pi(1-\lfloor q\rfloor/q)$. In this work, we choose
@@
When $q$ is noninteger, one may find that the Extended O(2) model with $h_{q}\rightarrow\infty$ and the Extended $q$-state clock model do not agree when $\beta\rightarrow 0$. For noninteger $q$, the $\mathbb{Z}_{q}$ symmetry is explicitly broken, and for example, the energy density of the Extended $q$-state clock model does not go to zero in the limit $\beta\rightarrow 0$ [7]. In contrast, the energy density in the Extended O(2) model does go to zero in the limit $\beta\rightarrow 0$ for all values of $h_{q}$. The discrepancy can be understood by considering the Boltzmann factor of the Extended O(2) model, which has the form